### v5

# Anomalous Andreev spectrum and transport in non-Hermitian Josephson junctions

Abstract
We propose a phase-biased non-Hermitian Josephson junction (NHJJ) composed of two superconductors mediated by a short non-Hermitian link. Such a NHJJ is described by an effective non-Hermitian Hamiltonian derived based on the Lindblad formalism in the weak coupling regime. By solving the Bogoliubov-de Gennes equation, we find that its Andreev spectrum as a function of phase difference exhibits Josephson gaps, i.e. finite phase windows with no Andreev (quasi-)bound states. The complex Andreev spectrum and the presence of Josephson gaps constitute particular spectral features of the NHJJ. Moreover, we propose complex supercurrents arising from inelastic Cooper pair tunneling to characterize the anomalous transport in the NHJJ. Additional numerical simulations complement our analytical predictions. We demonstrate that the Josephson effect is strongly affected by non-Hermitian physics.
I Introduction
Recently, there has been growing interest in non-Hermitian systems Bender (2007); El-Ganainy et al. (2018); Ashida et al. (2020); Bergholtz et al. (2021); Okuma and Sato (2023). Non-Hermitian physics emerges in a variety of fields such as disordered systems Zyuzin and Zyuzin (2018); Papaj et al. (2019), correlated solids Yoshida et al. (2019); Nagai et al. (2020), and open or driven systems Makris et al. (2008); Guo et al. (2009); Rotter (2009); Nakagawa et al. (2018); Xiao et al. (2019); Li et al. (2019); Zhang and Gong (2020). Different from the Hermitian case, the eigenvalues of non-Hermitian systems are complex in general. The complex-valued nature of the spectrum gives rise to unique features such as exceptional points Dembowski et al. (2001); Berry (2004); Kawabata et al. (2019a), point gaps Kawabata et al. (2019b); Zhou and Lee (2019), and enriched topological phases Lee (2016); Yao and Wang (2018); Kunst et al. (2018); Gong et al. (2018). Many of these particular phenomena have been theoretically studied and experimentally observed Lee (2016); Yao and Wang (2018); Kunst et al. (2018); Gong et al. (2018); Zhou and Lee (2019); Kawabata et al. (2019b, a); Zhang et al. (2020); Okuma et al. (2020); Yao et al. (2018); Zeng et al. (2016); San-Jose et al. (2016); Leykam et al. (2017); Shen et al. (2018); Yin et al. (2018); Yokomizo and Murakami (2019); Lee and Thomale (2019); Longhi (2019); Lee et al. (2019); Yamamoto et al. (2019); Avila et al. (2019); Borgnia et al. (2020); Budich and Bergholtz (2020); Yamamoto et al. (2021); Franca et al. (2022); Cayao and Black-Schaffer (2022); Jezequel and Delplace (2023); Qin et al. (2023); Guo et al. (2023); Cayao and Black-Schaffer (2023); Sun et al. (2023); Li et al. (2023); Kornich (2023); Zeuner et al. (2015); Ding et al. (2016); Xiao et al. (2020); Ozturk et al. (2021); Liang et al. (2022); Zhang et al. (2023); Liang et al. (2023); Xu et al. (2023). Whereas the research on characterizing novel non-Hermitian systems has achieved impressive progress, the impact of non-Hermiticity on quantum transport is largely unexplored despite some related efforts Zhu et al. (2016); Longhi (2017); Chen and Zhai (2018); Bergholtz and Budich (2019); Shobe et al. (2021); Michen and Budich (2022); Kornich and Trauzettel (2022); Sticlet et al. (2022); Geng et al. (2023); Isobe and Nagaosa (2023), especially for the cases involving superconductivity. For instance, an intriguing open question is how non-Hermiticity affects low-energy spectral features of superconducting systems and results in particular transport signatures.
The best settings to study superconducting transport are Josephson junctions (JJs). Generally, JJs consist of two superconductors separated by a (weak) link Likharev (1979); Beenakker (1992); Golubov et al. (2004); Tinkham (1996). Hermitian JJs have been extensively studied for decades. Phase-coherent quantum transport in Hermitian JJs gives rise to the dc Josephson effect: a phase-dependent supercurrent without voltage bias. In the short junction regime, the Josephson effect is entirely determined by the low-energy Andreev spectrum Furusaki (1999); Beenakker and van Houten (1991); Kwon et al. (2004); Fu and Kane (2009); Dolcini et al. (2015); Beenakker et al. (2013). When non-Hermiticity is introduced, we expect that the impact of non-Hermiticity on fundamental properties of JJs, such as the low-energy Andreev spectrum and the phase-biased supercurrent, is substantial. It is also of experimental relevance because JJs are always coupled to the environment and non-Hermitian parts of a Hamiltonian mimic certain aspects of this coupling.
In this paper, we investigate the low-energy spectrum and transport properties of a short non-Hermitian Josephson junction (NHJJ). The proposed NHJJ is built up of two superconductors connected by a non-Hermitian barrier [Fig. 1(a)]. We discover that the Andreev spectrum becomes complex-valued and exhibits Josephson gaps [purple regions in Fig. 1(b)]. Due to the complex-valued nature of the Andreev spectrum, the supercurrent becomes complex. Its imaginary part signifies the loss of quasiparticles due to inelastic Cooper pair tunneling. We also perform detailed tight-binding calculations to simulate experimental implementations of our proposal.
The paper is organized as follows. In Sec. II, we present the construction of the NHJJ and its effective non-Hermitian Hamiltonian from the Lindblad formalism in a weak coupling regime. In Sec. III, we derive the Andreev quasi-bound states and develop the concept of Josephson gaps. In Sec. IV, we present the inelastic Cooper pair tunneling and its consequent complex supercurrents. In Sec. V, we consider the experimental feasibility and numerical simulations of the NHJJ. Finally, we conclude in Sec.VI.
II Effective non-Hermitian Hamiltonian of the non-Hermitian Josephson junctions
We consider a minimal model of a one-dimensional (1D) NHJJ as sketched in Fig. 1(a). The junction is assumed to be coupled to a fermionic environment, constituting an open quantum system. We derive an effective NH Hamiltonian description for this system from the Lindblad formalism under certain approximations Breuer and Petruccione (2002); Daley (2014); Minganti et al. (2019); Manzano (2020); Roccati et al. (2022). The Hamiltonian for the total quantum system is composed of three parts
$H=H_{S}+H_{E}+H_{I},$ (1)
where $H_{S}$ is the system Hamiltonian, describing the JJ without coupling to the environment. The degrees of freedom in the environment are described by the Hamiltonian $H_{E}$. The interaction between system and environment is captured by $H_{I}$. The Hamiltonian of the system can be written in its eigenbasis as $H_{S}=\sum_{n_{s}}(\varepsilon_{n_{s}}-\mu_{s})C_{n_{s}}^{\dagger}C_{n_{s}}$ with eigenenergy $\varepsilon_{n_{s}}$ and eigenstate $|n_{s}\rangle$, where $C_{n_{s}}^{\dagger}$ ($C_{n_{s}}$) denotes creation (annihilation) operators of the system. They refer to the occupation of Andreev quasi-bound states. The fermionic environment is described by the Hamiltonian $H_{E}=\sum_{n_{e}}(\varepsilon_{n_{e}}-\mu_{e})C_{n_{e}}^{\dagger}C_{n_{e}}$. We assume that $\mu_{e}<\mu_{s}$, i.e. the chemical potential of the environment is smaller than that of the system. We envision that an external reservoir is coupled to the JJ via a quantum dot, cf. Fig. 1(a). The quantum dot works as a monitor of the tunneling particles between system and reservoir, as we elaborate on below. Note that the Hamiltonian $H_{E}$ describes the degrees of freedom of the external reservoir and the quantum dot. The coupling Hamiltonian is given by $H_{I}=\sum_{n_{e},n_{s}}V_{es}C_{n_{e}}^{\dagger}C_{n_{s}}+h.c.$. In general, the interaction term can be decomposed as $H_{I}=\sum_{i}S_{i}\otimes E_{i}$, where $S_{i}$ ($E_{i}$) is an operator acting on the system (the environment). We make a weak coupling assumption, in the sense, that the coupling between system and environment is weaker than the energy gap of Andreev levels.
The Lindblad master equation describes the dynamics of the system in the presence of dissipation due to the environment. The Lindblad master equation takes the form Breuer and Petruccione (2002); Daley (2014); Minganti et al. (2019); Manzano (2020); Roccati et al. (2022)
$\frac{d\rho_{S}}{dt}$ $=-i[H_{S},\rho_{S}]+\sum_{m}\gamma(\hat{L}_{m}\rho_{S}\hat{L}_{m}^{\dagger}-\frac{1}{2}\{\hat{L}_{m}^{\dagger}\hat{L}_{m},\rho_{S}\}),$ (2)
where $\rho_{S}$ is the reduced density matrix of the system and $\hat{L}_{m}$ the Lindblad operator describing the dissipative dynamics (including decoherence and loss processes) occurring at a positive rate $\gamma$ Daley (2014). We can rewrite Eq. (2) as
$\frac{d\rho_{S}}{dt}$ $=-i(H_{\mathrm{eff}}\rho_{S}-\rho_{S}H_{\mathrm{eff}}^{\dagger})+\sum_{m}\gamma\hat{L}_{m}\rho_{S}\hat{L}_{m}^{\dagger},$ (3)
where $H_{\mathrm{eff}}=H_{S}-\frac{i}{2}\sum_{m}\gamma\hat{L}_{m}^{\dagger}\hat{L}_{m}.$ Considering $\mu_{e}<\mu_{s}$, we assume that the energy modes of the environment coupled to the system are unoccupied Daley (2014). In this case, we have $\hat{L}_{m}=C_{m_{s}}$.
The dissipation term $-\frac{i}{2}\sum_{m}\gamma\hat{L}_{m}^{\dagger}\hat{L}_{m}$ describes the continuous loss of coherence and energy of the system to the environment. Without the quantum-jump term Yamamoto et al. (2019); Minganti et al. (2019); Roccati et al. (2022), we obtain the equation of motion $\frac{d\text{$\rho_{S}$}}{dt}=-i(H_{\mathrm{eff}}\rho_{S}-\rho_{S}H_{\mathrm{eff}}^{\dagger})$, which is equivalent to work with an effective non-Hermitian Hamiltonian
$H_{\mathrm{eff}}=H_{S}-\frac{i}{2}\sum_{m}\gamma\hat{L}_{m}^{\dagger}\hat{L}_{m}.$ (4)
The quantum-jump term $\sum_{m}\gamma\hat{L}_{m}\rho_{S}\hat{L}_{m}^{\dagger}$ is related to the abrupt change of quantum trajectories Minganti et al. (2019); Daley (2014). The Lindblad operator takes $\hat{L}_{m}=C_{m_{s}}$ in our case. Consider a state $|m_{s}\rangle=|0,\cdots,n_{m},\cdots\rangle$ in the system $H_{s}$. Then the quantum-jump term $\gamma C_{m_{s}}|m_{s}\rangle\langle m_{s}|C_{m_{s}}^{\dagger}$ indicates the annihilation of an excitation in the state $|m_{s}\rangle$. This excitation escapes into the environment. To monitor the quantum-jump process, we assume an auxiliary quantum dot between the junction and the external reservoir, as shown in Fig. 1(a). Note that this setup can be experimentally realized Deng et al. (2016). The quantum dot has a few energy levels. Hence, it is feasible to monitor its occupation e.g. by a nearby quantum point contact. By performing weak measurements and postselection techniques on the quantum dot Naghiloo et al. (2019), it is possible to only consider those trajectories where no quantum jumps happen during finite time windows. The average over many such quantum trajectories describes the loss physics due to the dissipation.
FIG. 1: (a) Sketch of a 1D NHJJ: two superconductors are connected by a short junction, which is coupled to the environment (a reservoir and a quantum dot (QD)). The QD is used to monitor the tunneling of quasiparticles between system and reservoir. The two superconductors possess a phase difference $\phi=\phi_{1}-\phi_{2}$. The wavy line indicates loss due to the coupling to the environment. (b) Complex Andreev spectrum as a function of $\phi$ within the superconducting gap $\Delta$ in the $s$-wave NHJJ. The purple regions indicate Josephson gaps. We choose $Z_{1}=0.6$ and $Z_{2}=0.2$ here. (c) The energy gap $E_{g}$ of Andreev spectrum as a function of the parameter $Z_{2}$ for different choices of $Z_{1}=0.4,0.5$ and $0.6$.
To capture this loss physics of the system in a minimal way, we employ a non-Hermitian potential at the junction as $-iV_{2}\delta(x)$ with $V_{2}>0$. This non-Hermitian potential mimics the loss of coherence and particles effectively. Furthermore, we assume that the junction between two superconductors also exhibits a real barrier potential described by $V_{1}\delta(x)$ with $V_{1}>0$. Hence, we take into account the following scattering potential at the junction
$U(x)=(V_{1}-iV_{2})\delta(x).$ (5)
This gives rise to the effective Hamiltonian $H_{\mathrm{eff}}=H_{S}+U(x)$.
Explicitly, the Bogoliubov-de Gennes (BdG) Hamiltonian to describe the NHJJ can be written as
$\mathcal{H}_{\mathrm{BdG}}(x)=\left(\begin{array}[]{cc}-\frac{\hbar^{2}\partial_{x}^{2}}{2m}-\mu+U(x)&\hat{\Delta}(x)\\ \hat{\Delta}^{\dagger}(x)&\frac{\hbar^{2}\partial_{x}^{2}}{2m}+\mu-U^{*}(x)\end{array}\right),$ (6)
where $m$ is the effective mass of the electrons, $\mu$ the chemical potential, and $\hat{\Delta}(x)$ the pairing potential. We argue that the particle-hole symmetry (PHS) of the type $U\mathcal{H}_{\mathrm{BdG}}^{*}U^{-1}=-\mathcal{H}_{\mathrm{BdG}}$ is physically relevant in our model (see Appendix A), where $U$ is a unitary matrix. Note that this particle-hole symmetry is labeled as $\mathrm{PHS}^{\dagger}$ in the literature Kawabata et al. (2019a, b). Considering the energy spectrum of the system, electron and hole excitations are described by the BdG equation
$\mathcal{H}_{\mathrm{BdG}}(x)\psi(x)=E\psi(x),$ (7)
where the eigenstate $\psi(x)=(u(x),v(x))^{T}$ is a mixture of electron and hole components, and $E$ is the excitation energy measured relative to the Fermi energy. The eigenenergy of $\mathcal{H}_{\mathrm{BdG}}$ always comes in pairs $E\longleftrightarrow-E^{*}$ under $\mathrm{PHS}^{\dagger}$.
III Andreev quasi-bound states and Josephson gaps
We consider an $s$-wave NHJJ, in which the two superconductors are both of $s$-wave type. Without loss of generality, we assume that left and right superconductors have a constant pairing potential of the same magnitude but different phases, i.e., $\hat{\Delta}(x<0)=\Delta$ and $\hat{\Delta}(x>0)=\Delta e^{i\phi}$ where $\phi$ is the phase difference across the junction. We are interested in the low-energy bound states existing within the superconducting gap. We obtain such Andreev quasi-bound states in the NHJJ by solving the BdG equation with proper boundary conditions. We may write the wave function of bound states as
$\psi_{B}(x)=\sum_{\eta,\alpha}A_{\eta\alpha}e^{i\alpha\sigma_{\eta}k_{\eta}x}\left(\begin{array}[]{c}e^{i\theta_{\eta\alpha}/2}\\ e^{-i\theta_{\eta\alpha}/2}\end{array}\right),$ (8)
where $\eta=e/h$, $\sigma_{e}\equiv 1,$ $\sigma_{h}\equiv-1$, and $\alpha=\pm$. The angles are defined as $\theta_{\eta+}\equiv\sigma_{\eta}\arccos(E_{B}/\Delta)+\phi$ and $\theta_{\eta-}\equiv\sigma_{\eta}\arccos(E_{B}/\Delta)$ with the energy of bound states $E_{B}$. $A_{\eta\alpha}$ is the coefficient of different evanescent modes. The bound states decay exponentially for $|x|\rightarrow\infty$ with a finite decay length $\lambda$. Therefore, the necessary condition for the existence of a bound state is given by $\mathrm{Re}(\lambda)>0$ with $\lambda=\frac{\hbar v_{F}}{\Delta}\frac{1}{\sqrt{1-E_{B}^{2}/\Delta^{2}}}$, where $v_{F}$ is the Fermi velocity.
III.1 Andreev quasi-bound states
By considering the boundary conditions of $\psi_{B}(x)$ at $x=0$, the secular equation to determine the energy of Andreev quasi-bound state is given by
$(Z^{2}+1)+2iZ_{2}\sqrt{\frac{\Delta^{2}}{E_{B}^{2}}-1}=\frac{\Delta^{2}}{E_{B}^{2}}\left(Z^{2}+\cos^{2}\frac{\phi}{2}\right).$ (9)
The parameter $Z_{j=1,2}$ for the strength of the non-Hermitian barrier potential is defined as $Z_{j}\equiv\frac{mV_{j}}{\hbar^{2}k_{F}}$, where $k_{F}$ is the Fermi wavevector and $Z^{2}\equiv\sum_{j}Z_{j}^{2}$.
In the clean case, $Z_{1}=Z_{2}=0$, the above equation reduces to the Kulik-Omel’yanchuk (KO) limit $E_{B}^{\pm}(\phi)=\pm\Delta\cos\frac{\phi}{2}$ Kulik and Omel’yanchuk (1978). For $Z_{1}\neq 0$ but $Z_{2}=0$, we obtain the known result $E_{B}^{\pm}(\phi)=\pm\Delta\sqrt{\frac{Z_{1}^{2}+\cos^{2}\frac{\phi}{2}}{Z_{1}^{2}+1}}$ Furusaki (1999). In this case, the Andreev spectrum has an energy gap $E_{g}=\frac{2\Delta Z_{1}}{\sqrt{Z_{1}^{2}+1}}$.
Even if the JJ couples weakly to the environment $V_{2}\ll E_{g}$, the imaginary term $2iZ_{2}\sqrt{\Delta^{2}/E_{B}^{2}-1}$ in Eq. (9) implies a complex Andreev spectrum, different from the Hermitian case Furusaki (1999); Beenakker and van Houten (1991); Kwon et al. (2004); Fu and Kane (2009); Dolcini et al. (2015); Beenakker et al. (2013). The dimensionless parameter $Z_{2}$ can be estimated as $Z_{2}\sim\frac{V_{2}L}{\hbar v_{F}}$, where $V_{2}$ describes the loss rate and $L$ is the junction length. Considering reasonable numbers, $Z_{2}$ is estimated to be of the order of $Z_{2}\sim 0.1$  [^1] If we assume $V_{2}\sim 1$ meV (a typical energy scale of our setup), junction length $L\sim 100$ nm, and Fermi velocity $v_{F}\sim 1\times 10^{6}$ m/s, the parameter $Z_{2}$ can be estimated as $Z_{2}\sim\frac{V_{2}L}{\hbar v_{F}}\approx 0.1$. The value of $Z_{2}$ can in principle be tuned by junction length $L$ and loss rate $V_{2}$ by changing the coupling between JJ and environment. . We focus on this experimentally relevant regime of $Z_{2}\ll 1$ below.
At special values $\phi=2n\pi$ with integer $n$, a trivial solution $E_{B}^{\pm}(\phi=2n\pi)=\pm\Delta$ exists, merging with the continuum. At $\phi=(2n+1)\pi$, we find that the Andreev spectrum becomes
$E_{B}^{+}[(2n+1)\pi]=\frac{Z^{2}\Delta}{\sqrt{Z^{4}+Z_{c}^{2}}},$ (10)
where $Z_{c}\equiv Z_{1}+iZ_{2}$ and $E_{B}^{-}=-(E_{B}^{+})^{*}$. The real energy gap $E_{g}$ obtained from Eq. (10) is illustrated in Fig. 1(c). Evidently, the energy gap $E_{g}$ is just slightly modified as we turn on a small $Z_{2}$ in the weak coupling regime.
For a general phase difference $\phi\in[\phi_{b},\phi_{t}]$, the secular equation can be written as
$\frac{E^{2}}{\Delta^{2}}=\frac{Z^{2}(Z^{2}-\sin^{2}\frac{\phi}{2})+\cos^{2}\frac{\phi}{2}+2Z_{1}^{2}\pm 2Z_{2}\sqrt{\cos^{2}\frac{\phi}{2}(Z^{2}-\sin^{2}\frac{\phi}{2})-Z_{1}^{2}}}{(Z^{2}+1)^{2}-4Z_{2}^{2}}.$ (11)
We plot the corresponding complex Andreev spectrum obtained from the above equation in Fig. 1(b). The real part of the complex Andreev spectrum indicates the physical energy while its imaginary part characterizes a finite lifetime of quasiparticles traveling across the junction. The phase boundaries are defined as $\phi_{b}\equiv 2n\pi+\phi_{0}(Z_{2})$ and $\phi_{t}\equiv(2n+1)\pi-\phi_{0}(Z_{2})$ with $\phi_{0}(Z_{2})\equiv 2\arcsin(\sqrt{2}Z_{2}/\sqrt{1+Z^{2}})$ obtained from the condition $\mathrm{Re}(\lambda)>0$. For $Z_{2}\ll 1$, this phase boundary can be approximated by $\phi_{0}(Z_{2})\simeq\frac{2\sqrt{2}Z_{2}}{\sqrt{1+Z_{1}^{2}}}$.
III.2 Josephson gaps
In general, the Andreev spectrum is a continuous function with respect to the phase difference $\phi$. However, we find that within a finite phase window $\Phi_{W}\equiv[2n\pi-\phi_{0}(Z_{2}),2n\pi+\phi_{0}(Z_{2})]$ with $n\in\mathbb{Z}$, low-energy Andreev quasi-bound states are not allowed at all within the superconducting gap $\Delta$. We call such phase windows Josephson gaps. Its gap value is $2\phi_{0}(Z_{2})=4\arcsin(\sqrt{2}Z_{2}/\sqrt{1+Z^{2}})$. The appearance of Josephson gaps can be understood as a consequence of the competition between phase-coherent transport of Cooper pairs across the junction and decoherence induced by the NH barrier. This competition can be understood by comparing the phase difference $\phi$ with the loss parameter $Z_{2}$. In the Hermitian case, due to phase coherence across the junction, quasiparticles move back and forth in the junction carrying the phase factor of superconductors imprinted by Andreev reflection Tinkham (1996); Nazarov and Blanter (2006); Golubov et al. (2004); Furusaki (1999). Andreev reflections lead to phase-coherent transport of Cooper pairs across the junction, i.e. dc supercurrent. However, the NH mechanism in the NHJJ gives rise to loss of coherence and inelastic Andreev reflections [see Fig. 2(a)] Nazarov and Blanter (2006); Daley (2014); Minganti et al. (2019). The magnitude of $Z_{2}$ determines the strength of decoherence. Hence, it is clear that the larger $Z_{2}$, the stronger is the coupling between system and environment. For phase differences $\phi\sim 0$ (mod $2\pi$) associated with a small supercurrent meaning a small number of transferred Cooper pairs, we observe Josephson gaps. This implies that phase-coherent transport of Cooper pairs is totally absent (thus no bound states). For larger supercurrents, i.e. larger phase differences, the pair breaking mechanism related to the barrier strength $Z_{2}$ is not sufficient to block the supercurrent. Nevertheless, decoherence modifies the supercurrent in this case. The Josephson-gap phase boundary $\phi_{0}(Z_{2})$ depends on the value of $Z_{2}$. In Fig. 2(b), we plot the Josephson-gap phase boundary as a function of $Z_{2}$. It grows almost linearly with increasing $Z_{2}$.
FIG. 2: (a) Sketch of inelastic Cooper pair tunneling. Transfer of Cooper pairs experiences possible decoherence and quasiparticles loss (wavy lines). (b) Josephson-gap phase boundary $\phi_{0}(Z_{2})$ as a function of $Z_{2}$. (c) Complex supercurrents as a function of $\phi$ with $Z_{1}=0.6$ and $Z_{2}=0.2$.
IV Inelastic Cooper pair tunneling and complex supercurrent
In (short) Hermitian JJs, the supercurrent originates from the coherent Cooper pair tunneling related to the Andreev spectrum Beenakker (1991, 1992); Furusaki (1999). In NHJJs, the Andreev spectrum becomes complex. The imaginary part of the Andreev spectrum indicates a finite lifetime of quasiparticles. Thus, the NH mechanism introduces loss of quasiparticles and decoherence during the Andreev reflection processes Breuer and Petruccione (2002); Nazarov and Blanter (2006); Daley (2014); Minganti et al. (2019) [see Fig. 2(a)], leading to inelastic Cooper pair tunneling  Devoret et al. (1990); Ingold and Nazarov (1994); Hofheinz et al. (2011); Westig et al. (2017); Grimm et al. (2019).
The supercurrent $I_{s}(\phi)$ can be obtained by taking the derivative of the free energy with respect to the phase difference $\phi$ Beenakker (1992). It leads to $I_{s}(\phi)=-\frac{2e}{\hbar}\sum_{n>0}\frac{dE_{n}(\phi)}{d\phi}$ with the Andreev spectrum $E_{n}(\phi)$ in Hermitian JJs Beenakker (1991, 1992). In NHJJ, the real part of the Andreev spectrum is the physical energy while the imaginary part corresponds to a finite lifetime of each eigenstate. The free energy in the ground state of NHJJs can be interpreted as the summation of real energy levels, broadened due to imaginary parts up to the Fermi energy Yamamoto et al. (2019). Following the same reasoning as in Hermitian JJs, we derive the supercurrent as
$I_{s}(\phi)=-\frac{2e}{\hbar}\frac{d\mathrm{Re}[E_{B}^{+}(\phi)]}{d\phi}+i\frac{2e}{\hbar}\frac{d\mathrm{Im}[E_{B}^{+}(\phi)]}{d\phi},$ (12)
where $E_{B}^{+}(\phi)$ is the complex Andreev spectrum. The real part $\mathrm{Re}[I_{s}(\phi)]$ represents the physical supercurrent flowing across JJs. The imaginary part $\mathrm{Im}[I_{s}(\phi)]$ stems from phase-dependent lifetimes of quasiparticles during inelastic Cooper pair tunneling. Thus, it is associated with the loss of quasiparticles from the JJ to the environment  [^2] The lifetime of quasiparticles becomes phase dependent. Thus, the magnitude of imaginary supercurrent indicates the response of quasiparticles loss as changing $\phi$.. Indeed, it is possible for either electrons or holes to escape the JJs after experiencing inelastic scattering at the NH barrier. If we consider scattering of electrons or holes, the scattering probability becomes $P(E,Z_{1},Z_{2})<1$, which means that some quasiparticles escape into the environment from the JJ (see Appendix B).
Explicitly for $s$-wave NHJJs, within the phase region $\phi\in[\phi_{b},\phi_{t}]$, the supercurrent as a function of $\phi$ is plotted in Fig. 2(c). Within the Josephson gaps, there is no supercurrent flowing across the junction. The supercurrent exhibits a jump at the phase boundary $\phi_{0}(Z_{2})$. We note that imaginary currents in normal states have been employed to describe delocalized behavior or dissipation of eigenstates Hatano and Nelson (1996); Zhang et al. (2022a); Li et al. (2021), while to the best of our knowledge, a complex supercurrent has not been addressed before.
V Experimental considerations and numerical simulations
Our analysis of NHJJs is experimentally relevant, considering the advanced fabrication techniques of JJs. The $s$-wave JJs could, for instance, be built by aluminum-based or InAs-based superconductor nanowires Doh et al. (2005); Chang et al. (2015); Hays et al. (2018). The local loss at the junction may stem from the coupling of JJs to the environment with a dissipative lead Zhang et al. (2022b); Liu et al. (2022). In the following, we provide more evidence from numerical simulations.
To perform tight-binding calculations, we discretize the Hamiltonian on a one-dimensional lattice by the long-wavelength approximation. The lattice constant is set at $a=1$. We divide the system into three parts: the left side and the right side are superconductors of length $L_{s}$, and the middle region is the normal part of length $L_{N}$. To mimic the loss at the junction, we set $L_{N}=1a$ and an on-site complex potential $V_{1}-iV_{2}$. The two superconductors have a paring potential $\Delta$ and a phase difference $\phi$.
FIG. 3: Tight-binding simulations of the $s$-wave NHJJ with a finite length. (a) Real part of energy spectra as a function of $\phi$. (b) Real supercurrent as a function of $\phi$ obtained from corresponding energy spectra in (a). We choose $V_{1}=0.8$, $V_{2}=0.3$, $\mu=1$ in units of $\Delta$, and $L_{s}=100a$ here.
Figure 3(a) shows the real part of the Andreev spectra as a function of $\phi$ for the NHJJ. Note the presence of the Josephson gap. Figure 3(b) illustrates the real part of the supercurrent obtained from the Andreev spectra corresponding to Fig. 3(a). The supercurrent vanishes within the Josephson gap.
To evaluate the phase boundary $\phi_{0}(Z_{2})$, we need the estimate $Z_{j}$ via $Z_{j}=\frac{mV_{j}}{\hbar^{2}k_{F}}$. The phase boundary is then obtained as $\phi_{0}(Z_{2})=2\arcsin\left(\frac{\sqrt{2}Z_{2}}{\sqrt{1+Z^{2}}}\right)$. For the parameters used in Fig. 3(b), we obtain $Z_{1}=0.566$, and $Z_{2}=0.212$, which gives $\phi_{0}(Z_{2})=0.165\pi$, in accordance with the plot.
VI discussion and conclusions
To summarize, we have studied the low-energy spectral features and transport properties of NHJJs. Our minimal model of NHJJs reveals anomalous NH physics emerging in JJs. We have introduced several conceptually new phenomena, including Josephson gaps and complex supercurrents, with no counterparts in the Hermitian case.
Our work extends the realm of the Josephson effect to NH physics and inspires potential applications: the NHJJ may function as a supercurrent switch by tuning the phase difference $\phi$ exploiting the presence of Josephson gaps. The impact of non-Hermiticity on other related phenomena, such as ac Josephson effects, Shapiro steps Tinkham (1996), JJ circuits Devoret and Schoelkopf (2013), and Josephson diode effects Ando et al. (2020), are interesting extensions of our work.
VII Acknowledgments
We thank Jan Budich, Haiping Hu, Jian Li, Zhuo-Yu Xian, and Chunxu Zhang for valuable discussions. This work was supported by the Würzburg-Dresden Cluster of Excellence ct.qmat, EXC2147, project-id 390858490, the DFG (SFB 1170), and the Bavarian Ministry of Economic Affairs, Regional Development and Energy within the High-Tech Agenda Project “Bausteine für das Quanten Computing auf Basis topologischer Materialen”.
Note added.—Recently, we came to notice a related work on NHJJs Cayao and Sato (2023).
Appendix A Relevant particle-hole symmetry of the model
In this section, we analyze the particle-hole symmetry that is physically relevant from a general form. From a field operator perspective, the particle-hole symmetry (PHS) mixes creation operator $\Psi^{\dagger}$ and annihilation operator $\Psi$ in a way
$C\Psi_{\alpha}C^{-1}=U_{\alpha\beta}\Psi_{\beta}^{\dagger},\ C\Psi_{\alpha}^{\dagger}C^{-1}=\Psi_{\alpha}U_{\alpha\beta}^{T},$ (13)
where $U$ represents a unitary matrix and $C$ denotes the PHS operator. A Hamiltonian $H=\Psi_{\alpha}^{\dagger}\mathcal{H}_{\alpha\beta}\Psi_{\beta}$ is particle-hole symmetric if $CHC^{-1}=H$. This implies that the first-quantized Hamiltonian transforms under PHS as
$C\mathcal{H}^{T}C^{-1}=-\mathcal{H},$ (14)
where $T$ is the transpose operation. In the Hermitian case, $\mathcal{H}^{T}=\mathcal{H}^{*}$. While in the non-Hermitian case, $\mathcal{H}^{T}\neq\mathcal{H}^{*}$ in general. Thus, the PHS comes in two kinds as Kawabata et al. (2019a):
$U\mathcal{H}^{*}U^{-1}$ $=-\mathcal{H},\ \text{$\mathrm{type}$\ I};$
$U\mathcal{H}^{T}U^{-1}$ $=-\mathcal{H},\ \mathrm{\text{$\mathrm{type}$\ I}I.}$ (15)
In the presence of PHS, the excitation energy of $\mathcal{H}$ always comes in pairs. For the two types of PHSs above, the corresponding energy pairs are $E\longleftrightarrow-E^{*}$ for type I PHS and $E\longleftrightarrow-E$ for type II PHS, respectively.
To determine which one of the two types of PHS on $\mathcal{H}$ is physically more relevant, we further consider the constraint of PHS on Green’s functions. On the one hand, the effective Hamiltonian $\mathcal{H}$ is directly related with the retarded Green’s function. Note that the poles of the retarded Green’s function yield the eigenvalues of the Hamiltonian, located in the lower half of the complex plane. On the other hand, the physical indication of the retarded Green’s function as a propagator is consistent with causality. The retarded Green’s function is defined in terms of field operators as $G_{\alpha\beta}^{R}(t)=-i\theta(t)\langle[\Psi_{\alpha}(t),\Psi_{\beta}^{\dagger}(0)]\rangle$, where $t$ is the time domain and $\theta(t)$ is a Heaviside step function. Considering PHS on field operators in Green’s function, it leads to
$G_{\alpha\beta}^{R}(t)$ $=-i\theta(t)\langle[\Psi_{\alpha}(t),\Psi_{\beta}^{\dagger}(0)]\rangle$
$=-i\theta(t)\langle C^{-1}C[\Psi_{\alpha}(t),\Psi_{\beta}^{\dagger}(0)]\rangle$
$=-i\theta(t)\langle C[\Psi_{\alpha}(t),\Psi_{\beta}^{\dagger}(0)]C^{-1}\rangle$
$=-U_{\alpha\gamma}^{*}\langle i\theta(t)[\Psi_{\gamma}^{\dagger}(0),\Psi_{\delta}(-t)]\rangle U_{\delta\beta}^{T}$
$=-U_{\alpha\gamma}^{*}G_{\gamma\delta}^{A}(-t)U_{\delta\beta}^{T}.$ (16)
By Fourier transformation and relating the retarded and advanced Green’s function by $[G_{\alpha\beta}^{R}(E)]^{*}=G_{\alpha\beta}^{A}(E)$, the retarded Green’s function fulfills
$U^{\dagger}[G^{R}(E)]^{*}U=-G^{R}(-E).$ (17)
This PHS constraint on retarded Green’s function is consistent with type I PHS acting on $\mathcal{H}$: The eigenvalues of the system (poles of the Green’s function) reside on the same side of the complex plane $\mathrm{Im}[E]\leq 0$. In contrast, from the constraint of type II PHS acting on $\mathcal{H}$, eigenvalues distribute symmetrically in the upper and lower half of the complex plane. This symmetric eigenenergy distribution is not consistent with the eigenvalues obtained from the retarded Green’s function. Thus, it is not compatible with the requirement of causality. Therefore, we argue that type I PHS acting on $\mathcal{H}$ is physically more relevant to describe non-Hermitian systems. This type I PHS is employed in the main text and labeled as $\mathrm{PHS^{\dagger}}$.
Appendix B Quasiparticle loss due to non-Hermitian barrier scattering
In this section, we obtain the reflection amplitudes (Andreev reflection and the normal reflection) and tunneling amplitudes (crossed Andreev tunneling and the normal tunneling) of quasiparticles after scattering at the non-Hermitian barrier. We show explicitly the loss of quasiparticles from the junction to the environment due to non-Hermitian scatterings. Consider an electron-like quasiparticle that is propagating from the left side of the junction. The wave function can be written as
$\psi(x)=\left(\begin{array}[]{c}u(x)\\ v(x)\end{array}\right)=\begin{cases}e^{ik_{e}x}\left(\begin{array}[]{c}u_{0}\\ v_{0}\end{array}\right)+A_{h-}e^{ik_{h}x}\left(\begin{array}[]{c}v_{0}\\ u_{0}\end{array}\right)+A_{e-}e^{-ik_{e}x}\left(\begin{array}[]{c}u_{0}\\ v_{0}\end{array}\right),&x<0;\\ A_{e+}e^{ik_{e}x}\left(\begin{array}[]{c}u_{0}e^{i\phi/2}\\ v_{0}e^{-i\phi/2}\end{array}\right)+A_{h+}e^{-ik_{h}x}\left(\begin{array}[]{c}v_{0}e^{i\phi/2}\\ u_{0}e^{-i\phi/2}\end{array}\right),&x>0.\end{cases}$ (18)
Note that $u_{0}^{2}=\frac{1}{2}\left(1+i\frac{\sqrt{\Delta^{2}-E^{2}}}{E}\right),\ v_{0}^{2}=\frac{1}{2}\left(1-i\frac{\sqrt{\Delta^{2}-E^{2}}}{E}\right)$. The wave vectors are defined as $\hbar k_{e}=\sqrt{2m(\mu+i\sqrt{\Delta^{2}-E^{2}})},\ k_{h}=k_{e}^{*}.$ The coefficients are determined by the boundary conditions:
$\psi(0+)$ $=\psi(0-),$ (19)
$\psi^{\prime}(0+)-\psi^{\prime}(0-)$ $=\frac{2m}{\hbar^{2}}(V_{1}\tau_{0}-iV_{2}\tau_{z})\psi(0+).$ (20)
By solving these linear equations, the four coefficients (Andreev reflection $A_{h-},$ normal reflection $A_{e-}$, normal tunneling $A_{e+}$, crossed Andreev tunneling $A_{h+}$) are obtained as
$A_{h-}$ $=-\frac{\Delta[E\sin^{2}\frac{\phi}{2}+\Omega(i\sin\frac{\phi}{2}\cos\frac{\phi}{2}-Z_{2})]}{(Z^{2}+1)E^{2}+2Z_{2}\Omega E-\Delta^{2}(Z^{2}+\cos^{2}\frac{\phi}{2})},$ (21)
$A_{e-}$ $=-\frac{Z^{2}\Omega^{2}-Z_{2}\Omega E-iZ_{1}\Omega^{2}}{(Z^{2}+1)E^{2}+2Z_{2}\Omega E-\Delta^{2}(Z^{2}+\cos^{2}\frac{\phi}{2})},$ (22)
$A_{e+}$ $=-\frac{\Omega E[Z_{2}\cos\frac{\phi}{2}+i(1+iZ_{1})\sin\frac{\phi}{2}]-\Omega^{2}[(1+iZ_{1})\cos\frac{\phi}{2}+iZ_{2}\sin\frac{\phi}{2}]}{(Z^{2}+1)E^{2}+2Z_{2}\Omega E-\Delta^{2}(Z^{2}+\cos^{2}\frac{\phi}{2})},$ (23)
$A_{h+}$ $=\frac{\Delta\Omega[Z_{2}\cos\frac{\phi}{2}-Z_{1}\sin\frac{\phi}{2}]}{(Z^{2}+1)E^{2}+2Z_{2}\Omega E-\Delta^{2}(Z^{2}+\cos^{2}\frac{\phi}{2})}.$ (24)
FIG. 4: Scattering probabilities as a function of $Z_{2}$ for fixed $Z_{1}=0.6$. We chose $\phi=\frac{\pi}{2}$ in this plot. We choose $E=1.5$ in unit of $\Delta$.
The parameter $\Omega$ is defined as $\Omega\equiv\sqrt{E^{2}-\Delta^{2}}$. The total scattering probability is defined as
$P(E,Z_{1},Z_{2})=\sum_{\eta=e/h,s=\pm}|A_{\eta s}|^{2}.$ (25)
In the Hermitian case, the probability is conserved at $P(E,Z_{1},Z_{2}=0)=1$. For nonzero $Z_{2}$, we find that the probability $P(E,Z_{1},Z_{2})<1$, which indicates the loss of quasiparticles from the junction to the environment [as shown in Figs. 4(a) and 4(b)]. The probability $P(E,Z_{1},Z_{2})$ decays continuously as increasing the strength of $Z_{2}$. Thus, the value of $Z_{2}$ measures the magnitude of loss.